SIPB IAP 2006 Feedback Form

We’d like to improve our IAP offerings and help you get the most out of your IAP. Please take a few minutes to fill out this form. You can either stick it into the envelope with everyone else’s now, or drop it in Interdepartmental or swing by our office (W20-557) and drop it off in person. (maybe something about how the survey is voluntary and confidential – how confidential do we want to guarantee?)

This class:

Please circle an answer. Feel free to elaborate. The pace of this class was: too slow – slow – just right – fast – too fast The amount of material covered was: far too little – not enough – just right – a bit much – felt like a 24-unit class The difficulty of this class was: too easy – easy – just right – hard – mind-blowingly difficult What did you like about the class? What would you like us to definitely keep?

What didn’t you like about this class? What do you want us to change?

Any other feedback on the class?

SIPB IAP 2006

How did you hear about this class? What could we have done to tell you about this and our other classes?

What other SIPB IAP classes have you taken this IAP, or want to take?

What classes would you like SIPB to offer next IAP?

Any comments on what you like or don’t like about what SIPB has done with IAP this year?

Thanks for helping us improve our IAP offerings to you! If you think of something else, please feel free to either email us at sipb-iap-sucker@mit.edu, or click the “Feedback” button on our webpage - http://sipb.mit.edu/iap. Thanks!

Mike Shaw, sipb-iap-sucker@mit.edu

SIPB, W20-557, x3-7788

http://sipb.mit.edu/iap